# ConRebSeg: A Segmentation Dataset for Reinforced Concrete Construction

### v1

@@ 6 Conclusion
In this work, we have presented a new dataset consisting of sequences of RGB images with segmentation labels for reinforcement concrete construction. The dataset is, as opposed to most datasets in construction, made publicly available to boost research in the area and contains 14,805 images in total, bigger than most task-specific datasets that focus on the construction sector. We have carried out an extensive analysis of our dataset and specifically analyzed variations in annotation styles. Furthermore, we have established and compared three baseline segmentation models: U-Net, DeepLabV3, and YOLOv8L-seg. Our study revealed that the YOLOv8L-seg model trained from pre-trained weights is superior in performance compared to U-Net and DeepLabV3. Moreover, we have explored the effects of withholding training data while maintaining an equal class distribution and have found that the performance of the model increases with more available training data. In addition, we have explored the influence of different annotation styles on the model performance by removing anomalous annotations from the training data and have found no significant change. Lastly, our final evaluation showed that the model is precise in predicting areas with exposed reinforcement bars, but fails to generalize and recognize these areas in scenes that deviate from those depicted in the train split.
Our study concludes that there is still a significant lack of CV data for construction tasks. Unlike other sectors, where digitization has been adopted swiftly in the past decades, construction has been traditionally slow to adopt digital technologies. This has resulted in low data availability and thus, a challenging environment to develop deep learning-based solutions to be used in autonomous robotic systems in construction. These robotic systems will be a necessity in the future, as the construction sector is facing labor shortages in the upcoming years. We think the community will benefit from more publicly available datasets and a common data lake, where researchers will upload data along with annotations, even if annotations styles within a category are vastly different as our study showed the minimal impact of this phenomenon.
@@ Acknowledgements
This work has been funded and supported by the EU Horizon Europe project “RobetArme” under the Grant Agreement 101058731. We extend our gratitude to Christiansen & Essenbaek A/S (CEAS) and Markus Schmidtke, for sharing their domain expertise and organizing access to their construction sites for data-capturing purposes.
@@
The self-collected sequences of the dataset can be found on DTU Data [50]. The code to initiate the dataset including the dataset management tool can be found on GitHub in our repository [^2] https://github.com/DTU-PAS/ConRebSeg.
@@
Lazaros Nalpantidis: Funding acquisition, Supervision, Writing - Review and Editing

### v2

@@ 7 Conclusion
This paper presented a dataset consisting of temporally ordered RGB images with segmentation labels for reinforced concrete construction. The dataset is, as opposed to most datasets in construction, made publicly available to boost research in the area and contains 14,805 images in total, bigger than most task-specific datasets that focus on the construction sector. We have carried out an extensive analysis of our dataset and specifically analyzed variations in annotation styles. Furthermore, we have established and compared three baseline segmentation models: U-Net, DeepLabV3, and YOLOv8L-seg. Our study revealed that the YOLOv8L-seg model trained from pre-trained weights is superior in performance compared to U-Net and DeepLabV3. Moreover, we have explored the effects of withholding training data while maintaining an equal class distribution and have found that the performance of the model increases with more available training data. In addition, we have explored the influence of different annotation styles on the model performance by removing anomalous annotations from the training data and have found no significant change. In addition, we implemented the model on an embedded GPU and found a good compromise between model accuracy and frame rate using FP16 precision. Lastly, our final evaluation showed that the model is precise in predicting areas with exposed reinforcement bars, but fails to generalize and recognize these areas in scenes that deviate from those depicted in the train split.
Our study concludes that there is still a significant lack of CV data for construction tasks. Unlike other sectors, where digitization has been adopted swiftly in the past decades, construction has been traditionally slow to adopt digital technologies. This has resulted in low data availability and, thus, a challenging environment to develop deep learning-based solutions to be used in autonomous robotic systems in construction. These robotic systems will be a necessity in the future, as the construction sector is facing labor shortages in the upcoming years. We think the community will benefit from more publicly available datasets and a common data lake, where researchers will upload data along with annotations, even if annotations styles within a category are vastly different, as our study showed the minimal impact of this phenomenon.
@@ Acknowledgements
This work has been funded and supported by the EU Horizon Europe project “RobétArmé” under the Grant Agreement 101058731. We extend our gratitude to Christiansen & Essenbaek A/S (CEAS) and Markus Schmidtke, for sharing their domain expertise and organizing access to their construction sites for data-capturing purposes.
@@
The self-collected scenes of the dataset can be found on DTU Data [56]. The code to initiate the dataset including the dataset management tool can be found on GitHub in our repository [^3] https://github.com/DTU-PAS/ConRebSeg.
@@
Lazaros Nalpantidis: Funding acquisition, Supervision, Writing – Review and Editing